OpenStudy (anonymous):

One litter of paint is needed to cover all 6 sides of a cubical block. How many liters will be needed to cover all 6 sides of a second cubical block whose edge is twice as long as an edge of the first block? (a) 2 (b) 4 (c) 6 (d) 8 The second hand of a clock is 5 cm long. In one hour, the tip of the second hand travels a distance of (a) 600(pi)cm (b) 3600(pi)cm (c) 6000(pi)cm (d)36000(pi)cm Is the answer for the second question B?

OpenStudy (anonymous):

I'm going to focus on the second question since you've gotten an answer for it. In one minute, the second hand travels one circumference (once around the clock) and in one hour, it'll travel 60 circumferences. One circumference is equal to pi times the diameter. The diameter is 10 cm (twice the radius). So your answer would be: 60(10)(pi) = ?

OpenStudy (tkhunny):

Consider the first cube. Edge is length E. Total Surface Area is \(6\cdot E^{2}\) Now, double it: \(6\cdot (2E)^{2}\). How is that different? A ratio might help.

OpenStudy (anonymous):

219.8

OpenStudy (tkhunny):

What is that? Try this: \(\dfrac{6\cdot (2E)^{2}}{6\cdot E^{2}} = 8\) Does that mean anything, yet?

OpenStudy (anonymous):

nvm about that. It's an..equation..

OpenStudy (tkhunny):

Have you seen choice D in the first question?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

For the first question, assume that the length of a side of the small cube is x. So, a face is x square. Since you have six faces, the total surface area is 6x squared. Because the small edge's length is x, the larger cube's edge will be 2x. So, a face is (2x) squared, or 4x squared. Since there are still six faces, the total surface area will be 24x squared. In order to cover a small cube completely (6x squared), you need one liter of paint, so it you have a large cube (24x squared), you will need 4 times as much paint, or 4 liters of paint

OpenStudy (anonymous):

i hope that helped... a lot of information thought. basically, you need 4 liters of paint

OpenStudy (anonymous):

also, the answer to the second one is "a"

OpenStudy (anonymous):

Okay! Thank you! It helped a lot btw~

OpenStudy (tkhunny):

Right. \(2^{2} = 4 \ne 8\). Sorry about that.